OpenStudy (abmon98):

a geometric series has the following properties; the 1st and 2nd have a sum of -4 and the 4th and 5th term have a sum of 108. Find the 1st term and the common ratio of the series. Explain why the series has no sum to infinity.

OpenStudy (campbell_st):

so if you look at the formula for each term you get a term in a geometric series is \[a_{n} = ar^{n -1}\] a = 1st term and r is the common ratio 1st and 2nd sum equation 1 \[a + ar = -4\] 4th and 5th sum equation 2 \[ar^3 + ar^4 = 108\] with a little factoring of the equation 2 you get \[r^3(a + ar) = 108\] now substitute equation 1 \[r^3 (-4) = 108\] now you can solve for r. when you get r, substitute it into the 1st equation to find a hope it helpss

OpenStudy (anonymous):

a, ar, ar^2, ar^3 and ar^4 are the first 5 terms a + ar = -4 ar^3 + ar^4 = 108 solve a = 2 and r = -3

OpenStudy (abmon98):

Thank you so much for your help sourwing and campbell_st